OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 2

OpenStudy (anonymous):

A. (y + 2) + y = 6 B. x + (y + 2) = 6 C. x = (x – 6) + 2 D. x + y = y + 2

OpenStudy (anonymous):

@igreen

Nnesha (nnesha):

substitution method to replace something with variable

Nnesha (nnesha):

x + y = 6 x = y + 2 so here  so here is x by itself